how much faster does a c3 alkane (propane) diffuse than 233ThF6 (g)
Answer #1

Graham's Law of Diffusion:

(Rate1)/(Rate2) = Mass2/Mass1

The mass of propane gas (mass 1) = 44.1 g/mol.
The other gas 233ThF6 has a molecular mass (mass 2) = 233 + 19x6 = 347
and we will assign its rate the value of 1 (rate 2).
Now we solve for Rate 1.
From Graham's Law,:
(Rate 1)² = (Rate 2)² × Mass 2 ÷ Mass 1
plug in the numbers:
(Rate 1)² = 1² × 347 ÷ 44.1
(Rate 1)²= 7.87
Rate 1 = 2.8

Propane is 2.8 times faster diffuses than 233ThF6 (g).